An isoceles triangle has two side lengths of 9x+3. The perimeter of the triangle is 30x+10.
a) Determine the ratio of the base to the perimeter, in simplified form. State the restriction.

Answer by stanbon(75887)

An isoceles triangle has two side lengths of 9x+3. The perimeter of the triangle is 30x+10.
a) Determine the ratio of the base to the perimeter, in simplified form. State the restriction.
base = 30x+10 - 2(9x+3) = 30x-18x + 10-6 = 12x+4
--------------------------
Ratio of base to perimeter = (12x+4)/(30x+10) = [4(3x+1)/(10(3x+1)] = 4/10 = 2/5
-------------------
Restriction:: 3x+1 cannot be zero ; x cannot be -1/3.
--------------------
Cheers,
Stan H.
------------------